OpenStudy (maddie82502):

3x+1/14=2k/8 fan and medal if help

OpenStudy (anonymous):

Is this correct? \[3x+\frac{1}{14}=\frac{2k}{8}\]

OpenStudy (maddie82502):

I think

OpenStudy (anonymous):

Okay, what would you do first? We need to get k by itself.

OpenStudy (maddie82502):

divide

OpenStudy (anonymous):

Wait sorry I've just re-read the equation. What exactly are you trying to do in this question? There are two variables.

OpenStudy (maddie82502):

solve for x

OpenStudy (maddie82502):

@Hero @Brooklyn_Boiii @shishi990 @tkhunny

OpenStudy (anonymous):

Okay, so we need to get x by itself, not k.\[3x+\frac{1}{14}=\frac{2k}{8}\] What would you do first?

OpenStudy (maddie82502):

so do we add the 3 to the 1/14

OpenStudy (anonymous):

Nope, because it's 3x we can't just add it. We want to move things over to the right to get the 3x by itself, then we can divide both sides by 3 to leave us with x=...

OpenStudy (maddie82502):

ohh ok dokie